# Semantics-Preserving Evasion of LLM Vulnerability Detectors

### v1

@@ 4.1 Datasets Construction
Unified benchmark and design goals. To isolate robustness to semantics-preserving edits from confounders such as duplication leakage and length artifacts, we construct a single unified benchmark, Unified-VUL-N (N=5000N{=}5000). We pool vulnerable functions from PrimeVul, BigVul, and DiverseVul, enforcing three strict controls: (i) global cross-source de-duplication to prevent inflation from repeated functions, (ii) a fixed source composition to standardize evaluation across models, and (iii) a shared prompt budget to ensure valid comparisons.
Deduplication by normalization-only hashing. To prevent evaluation inflation from surface-level duplicates, we canonicalize code logic (stripping comments and formatting) to define a deterministic equivalence class via SHA-256 hashing. This normalization is used strictly for cross-source auditing; models are always evaluated on the original, unmodified source code. Detailed normalization rules are provided in Appendix B.1.
Filtering and Sampling. We restrict code to $\leq 4096$ tokens to reserve headroom for adversarial carriers and eliminate truncation-driven artifacts. A deterministic, quota-controlled sampling strategy ensures a stable and auditable benchmark composition (also see Appendix B.1).
@@ 4.2 Evaluation Protocol
Evaluation Decomposition: Conditional Evasion. To isolate robustness from clean-accuracy capabilities, we evaluate attacks conditionally on Clean True Positives ($\mathrm{TP}_{\mathrm{clean}}$). Since a pre-existing error cannot be ”further evaded,” applying attacks to the entire dataset would inflate perceived robustness. We define $\mathrm{Flip}(A)$ as the subset of $\mathrm{TP}_{\mathrm{clean}}$ successfully evaded under attack $A$, and $\mathrm{Resist}(A)$ as those that maintain detection.
Primary Robustness Metrics. We report three metrics to characterize the Robustness & Accuracy Gap:
@@
    Complete Resistance (CR). The fraction of vulnerabilities that withstand all evaluated attack families ($\cap_{k}\mathrm{Resist}(A_{k})/|\mathrm{TP}_{\mathrm{clean}}|$). This serves as the security floor, indicating the proportion of reliable detections that are truly invariant to semantic perturbations.
@@
    End-to-End Recall Drop ($\Delta\mathrm{TPR}$). The absolute decrease in detection coverage on the full dataset ($|\mathrm{Flip}_{\cup}|/N$). This measures the operational risk of deploying the model in an adversarial environment.
Formal derivations and single-family variants are detailed in Appendix B.2.
@@
Table 1: Baseline clean coverage on the unified C/C++ evaluation set. Absolute counts and $\mathrm{TPR}_{\mathrm{clean}}$ are reported; attack metrics ($|\mathbf{Flip}_{\mathrm{union}}|$, $|\mathbf{Resist}_{\mathrm{complete}}|$) appear in §5.1.
Model NN $|\mathbf{TP}_{\mathrm{clean}}|$ $|\mathbf{FN}_{\mathrm{clean}}|$ $\mathrm{TPR}_{\mathrm{clean}}$ (%)
Qwen2.5-Coder-32B 5000 1121 3879 22.42
Llama3.1-8B 5000 1552 3448 31.04
CodeAstra 5000 3454 1546 69.08
GPT-4o 5000 2300 2700 46.00
GPT-5-mini 5000 3681 1319 73.62
@@ 5.1 Main Results: Overall Attack Effectiveness
One would expect that a detector achieving 70% clean accuracy provides reliable security coverage for the vulnerabilities it identifies. Our evaluation reveals the opposite: across all models except GPT-5-mini, over 87% of correctly detected vulnerabilities can be flipped to false negatives using only semantics-preserving edits (Figure 2). CodeAstra, despite correctly flagging 3,454 vulnerabilities on clean inputs, maintains robust detection for only 4 functions under on-target attacks (CR = 0.12%), a 99.88% collapse in detection integrity.
This pattern occurs under realistic access assumptions: surrogate-transfer attacks on black-box APIs (GPT-4o, GPT-5-mini) and operational black-box models (Qwen2.5-Coder-32B), supplemented by on-target optimization for accessible models (Llama3.1-8B, CodeAstra) to establish upper bounds. Across this threat landscape, GPT-5-mini is an exception, exhibiting substantially lower attack success (44.04% union $\mathrm{ASR}_{\mathrm{cond}}$) compared to other models (87%–100%), suggesting stronger invariance to carrier-based perturbations. These results indicate that clean benchmark accuracy is not a security guarantee. An attacker needs only to evade the subset of cases a detector would otherwise catch, and our results show this is achievable through innocuous transformations like identifier renaming or comment insertion (Table 2; clean baseline in Table 1).
Figure 2: Overall attack effectiveness across models. Solid bars: transfer attacks using universal strings from Qwen2.5-Coder-14B surrogate. Hatched bars: on-target optimization (Llama3.1-8B, CodeAstra only). (a) Union $\mathrm{ASR}_{\mathrm{cond}}$ across carriers. (b) End-to-end recall decrease ($\Delta\mathrm{TPR}$).
Benign-class invariance. To verify that our edits induce targeted false negatives rather than indiscriminate label corruption, we evaluate attacks on a held-out set of 500 ground-truth benign functions. Across all models and attack methods, we observe 100% benign-class invariance: every function predicted as BENIGN under clean inference remains BENIGN after applying our transformations. This supports that the observed flips reflect targeted integrity violations rather than indiscriminate label corruption.
@@ 5.2 Carrier-Specific Vulnerabilities: Preprocessor Attacks Dominate
We break down attack effectiveness by carrier under surrogate transfer (Table 2). Three patterns emerge. First, no single carrier dominates: ifdef achieves near-complete evasion on Llama3.1-8B (99.03%) and Qwen2.5-Coder-32B (98.84%) but achieves moderate success elsewhere. Second, preprocessor carriers often outperform identifier substitution, suggesting over-reliance on non-executable boundary tokens. Third, GPT-5-mini exhibits uniformly low ASR (10–28%), indicating broader invariance. Besides, Dead-branch carriers are also effective on several models, indicating limited dead code alignment.
Model-specific patterns. Most models show high vulnerability to preprocessor carriers, with Llama3.1-8B and Qwen2.5-Coder-32B near-ceiling (85–99%). CodeAstra presents a distinct profile: identifier substitution transfers poorly (4.72%), while boundary carriers remain effective (42–79%). On-target optimization reverses this pattern (§5.4).
@@
5.3 Sanitization Eliminates Carriers but Induces Prediction Drift
@@
Qwen2.5-Coder-32B 91.53→0% 98.84→0% 85.81→0% +5.2%
Llama3.1-8B 64.50→0% 99.03→0% 96.33→0% +8.6%
CodeAstra 78.60→0% 42.59→0% 75.48→0% +0.1%
Given the high success of comment- and preprocessor-based carriers, we evaluate a simple input-sanitization baseline that removes these carrier surfaces prior to inference. The pipeline (i) strips C/C++ line and block comments, and (ii) canonicalizes/removes preprocessor directives before passing code to the detector. We treat this as a carrier-elimination baseline rather than a semantics-aware defense.
Table 3 shows that sanitization successfully drives ASR to zero for targeted carriers. However, it induces substantial prediction drift: stripping benign comments and preprocessor directives from unperturbed, ground-truth vulnerable functions increases clean detection rates by up to 8.6% (Llama3.1-8B: 1552→1685 true positives). This reveals that models utilize non-executable lexical context as decision-relevant features, indicating a lack of semantic invariance even absent adversarial perturbations (detailed drift analysis in Table 10, Appendix D).
Moreover, sanitization is not a free lunch: comments and preprocessor directives appear in 57.0% and 12.8% of functions, respectively, meaning this defense rewrites a substantial fraction of inputs. We exclude dead-code elimination, as it risks mis-eliminating live code without compiler-level context (Tu et al., 2024). Furthermore, sanitization fails to address executable-looking carriers like identifier substitution and dead-branch insertions.
@@ 5.4 Transfer Attacks Succeed on Black-Box APIs; On-Target Reveals Complex Interactions
Table 2 presents transfer results where universal strings optimized on Qwen2.5-Coder-14B are applied unchanged to all targets. This models a low-barrier adversary: optimization is performed once, and resulting strings are reused across vulnerable functions and deployed detectors. Across black-box and operational black-box targets, transfer achieves 44.04%–99.03% union $\mathrm{ASR}_{\mathrm{cond}}$, establishing that API opacity alone does not prevent evasion under semantics-preserving perturbations. However, transferability is not uniform: GPT-5-mini exhibits materially lower success (44.04%) than other models (87–99%), indicating that robustness varies across model families even under identical carriers.
@@
On-target optimization reveals divergent vulnerability profiles. While transfer attacks establish a practical lower bound on exploitability, they may underestimate the true vulnerability of a model. To quantify the gap between low-cost transfer and worst-case security risk, we perform on-target universal GCG for Llama3.1-8B and CodeAstra (accessible models). Table 4 compares transfer and on-target regimes across carriers.
Model Regime $\mathrm{ASR}_{\mathrm{cond}}$ (%)
Idsub Comment ifdef macro Dead-br
Llama3.1-8B Transfer 34.66 64.50 99.03 96.33 59.79
On-target 43.49 42.01 99.03 65.46 86.66
CodeAstra Transfer 4.72 78.60 42.59 75.48 72.12
On-target 85.41 63.06 16.62 98.87 97.51
@@
For Llama3.1-8B, on-target optimization improves identifier substitution (34.66% $\rightarrow$ 43.49%) and dead-branch insertion (59.79% $\rightarrow$ 86.66%), but degrades comment (64.50% $\rightarrow$ 42.01%) and macro (96.33% $\rightarrow$ 65.46%). This suggests that universal strings optimized on a surrogate can outperform target-specific optimization for certain carriers, likely because surrogate-optimized universal strings exploit more transferable boundary-token sensitivities.
CodeAstra exhibits a clear asymmetry. Transfer achieves 4.72% identifier success but 42.59% for ifdef. On-target optimization inverts this: identifier reaches (85.41%) while ifdef collapses (16.62%).
@@
We hypothesize two mechanisms. For identifier substitution, CodeAstra’s loss landscape provides smooth gradients: GCG efficiently discovers variable-name perturbations that flip predictions. For ifdef, on-target GCG likely encounters optimization difficulty, so the optimizer converges to poor local minima that fail to exploit the vulnerabilities that surrogate-optimized strings (accidentally) discovered through multi-model training. This suggests that gradient-based optimization success depends not only on model vulnerabilities but also on the optimizability of the attack surface: some carrier spaces may be inherently harder to optimize even when vulnerabilities exist.
@@
Qwen2.5-Coder-32B: Surrogate capacity and on-target upper bound. For Qwen2.5-Coder-32B (operational black-box), we ablate identifier substitution across surrogate capacities and on-target optimization (Table 5). Transferability is non-zero even for small surrogates (0.5B/1.5B:  6-7%), but increases with capacity (14B: 26.94%), indicating a representation-capacity threshold. On-target universal GCG achieves 92.50%, providing a clear upper bound when gradients are available. This gap (26.94% $\rightarrow$ 92.50%) quantifies the security improvement from restricting adversarial access: surrogate transfer yields a conservative lower bound under practical compute constraints, while on-target optimization demonstrates worst-case exploitability.
On-target universal GCG (Qwen2.5-Coder-32B) 92.50 1038 / 1121
Implications. These findings highlight three key insights. First, transferability amplifies risk by enabling low-cost, reusable attacks against multiple deployed detectors, including API-access models. Second, on-target optimization does not uniformly strengthen attacks- CodeAstra’s inverted profile and Llama’s macro degradation suggest complex cross-carrier interactions that warrant further investigation. Third, the gap between transfer and on-target regimes varies by model and carrier (4.72% $\rightarrow$ 85.41% for CodeAstra identifier vs. 99.03% $\rightarrow$ 99.03% for Llama ifdef), indicating that vulnerability assessment requires evaluating both threat models.
@@
Figure 3: Robustness against Non-GCG Baselines. We compare prompt injection against random_idsub. While prompt injection achieves high success on some models (e.g., Llama3.1, CodeAstra), random perturbations ($\sim$10%) remain far less effective than our optimized attacks, confirming that code syntax requires targeted optimization rather than random noise.

### v2

@@ 4.1 Datasets Construction
Unified benchmark. We construct Unified-VUL-N (N=5000N{=}5000) by pooling vulnerable C/C++ functions from PrimeVul, BigVul, and DiverseVul with three controls: global cross-source SHA-256 deduplication (on comment-stripped, whitespace-normalized code), a fixed per-source composition (PrimeVul 3000, BigVul 1000, DiverseVul 1000), and a shared prompt budget (${\leq}4096$ tokens via Qwen2.5-Coder-14B tokenizer). Deduplication targets cross-source inflation; models are always evaluated on the original, unmodified source code. Full normalization rules and sampling procedure are provided in Appendix B.1.
@@ 4.2 Evaluation Protocol
Conditional Evasion. We evaluate attacks on Clean True Positives ($\mathrm{TP}_{\mathrm{clean}}$) only: $\mathrm{Flip}(A)$ is the subset flipped under attack $A$; $\mathrm{Resist}(A)$ is the subset that maintains detection. Applying attacks to pre-existing misses would inflate apparent robustness.
Robustness Metrics. We report three metrics to characterize the Robustness & Accuracy Gap:
@@
    Complete Resistance ($\mathrm{CR}$). The fraction of vulnerabilities that resist all evaluated attack variants: $\mathrm{CR}={\bigl|\bigcap_{k\in\mathcal{K}}\mathrm{Resist}(A_{k})\bigr|}\big/{|\mathrm{TP}_{\mathrm{clean}}|}$ where $\mathcal{K}=\{\text{idsub, comment, macro, {\#ifdef}, dead-branch}\}$. Each $A_{k}$ uses one canonical insertion site per variant (leading comment header, macro-name, #ifdef DEBUG_MODE, dead-branch before final brace; ablations over placements are in Table 6). CR is a lower bound on the security floor under the evaluated threat set: a vulnerability counted in CR resists every canonical attack; a vulnerability not in CR can be evaded by at least one. Because CR is defined relative to $\mathcal{K}$, it can only decrease as $\mathcal{K}$ is extended (e.g., adding more carrier positions or surrogate models).
@@
    End-to-End Recall Drop ($\Delta\mathrm{TPR}$). The absolute decrease in detection coverage on the full dataset ($|\mathrm{Flip}_{\cup}|/N$). This measures the risk of deploying the model in an adversarial environment.
Formal derivations and single-family variants are detailed in Appendix B.2. For random_idsub, we report mean $\pm$ std over K=3K{=}3 random seeds. GCG-based attacks use a fixed seed and support set $\mathcal{X}_{\text{opt}}$; point estimates reflect deterministic optimization under fixed conditions.
Statistical qualification. Multi-seed analysis (on-target identifier substitution, N=3N{=}3 support sets, seeds 1/21/42) yields $\mathrm{ASR}_{\mathrm{cond}}$: CodeAstra $80.7\%{\pm}9.1$, Llama3.1-8B $31.1\%{\pm}11.3$, StarCoder2-15B $17.6\%{\pm}6.8$. All seeds produce successful attacks, confirming systematic exploitability; spread (${\leq}12$ pp) reflects support-set variability. Other model/carrier values are single-run estimates under seed 21 with fixed $\mathcal{X}_{\text{opt}}$.
@@ 5.1 Main Results: Overall Attack Effectiveness
One would expect that a detector achieving 70% clean accuracy provides reliable security coverage for the vulnerabilities it identifies. Our evaluation reveals the opposite: across all models except GPT-5-mini, over 87% of correctly detected vulnerabilities can be flipped to false negatives using only syntax-preserving edits (Tree-sitter-validated; compilation-preserving for comment, preprocessor, and dead-branch carriers and 90.9%-preserving for identifier substitution on the compilable subset; Table 1). CodeAstra, despite correctly flagging 3,454 vulnerabilities on clean inputs, maintains detection for only 4 functions under on-target attacks (CR = 0.12%), a 99.88% reduction in integrity.
This pattern occurs under realistic access assumptions: surrogate-transfer attacks on black-box APIs (GPT-4o, GPT-5-mini) and cost-limited gray-box models (Qwen2.5-Coder-32B), supplemented by on-target optimization for white-box models (Llama3.1-8B, CodeAstra, StarCoder2-15B). Across this threat landscape, GPT-5-mini is a notable exception, exhibiting substantially lower attack success (44.04% union $\mathrm{ASR}_{\mathrm{cond}}$) compared to other models (87%–100%), suggesting broader invariance.
Table 1 gives end-to-end post-attack metrics. Clean benchmark recall is not a security guarantee: an attacker needs only to evade the cases a detector correctly identifies, and our results show this is achievable via innocuous transformations (full carrier breakdown in Table 2).
Table 1: Clean coverage and post-attack robustness (N=5000N{=}5000). Transfer: universal strings from Qwen2.5-Coder-14B surrogate applied without modification. On-target: GCG optimized directly on the target model. Models with only one row were evaluated under that regime only.
Model Regime $\mathrm{TPR}_{\mathrm{clean}}$ (%) Union $\mathrm{ASR}_{\mathrm{cond}}$ (%) $\mathrm{CR}$ (%) $\Delta\mathrm{TPR}$ ($p$)
Llama3.1-8B Transfer 31.04 99.68 0.32 30.94
On-target 99.87 0.13 30.99
CodeAstra Transfer 69.08 95.00 5.00 65.63
On-target 99.88 0.12 69.01
StarCoder2-15B Transfer 99.04 87.72 12.28 86.88
On-target 90.39 9.61 89.52
Qwen2.5-Coder-32B Transfer 22.42 99.38 0.62 22.28
GPT-4o Transfer 46.00 87.00 13.00 40.02
GPT-5-mini Transfer 73.62 44.04 55.96 32.42
Figure 2: Non-GCG baselines. prompt injection tests output-format harness vulnerability (outside threat model); random_idsub tests code-understanding robustness (within threat model). Random Token Insertion on comment carriers achieves 13.2% vs. GCG’s 91.5%.
Benign-class invariance. To verify that our edits induce targeted false negatives rather than label corruption, we evaluate attacks on a held-out set of 500 ground-truth benign functions. Across all models and attack methods, we observe 100% benign-class invariance: every function predicted as BENIGN under clean inference remains BENIGN after applying our transformations. This supports that the observed flips reflect targeted integrity violations rather than indiscriminate label corruption.
@@ 5.2 Carrier-Specific Vulnerabilities: Preprocessor Attacks Dominate
StarCoder2-15B 1.31% 4.00% 0.18% 86.83% 1.58%
GPT-4o 31.09% 65.48% 71.61% 42.35% 19.09%
GPT-5-mini 10.62% 18.34% 28.14% 21.57% 17.58%
Table 2 breaks down attack effectiveness by carrier under surrogate transfer. No single carrier dominates across models: ifdef achieves near-complete evasion on Llama3.1-8B (99.03%) and Qwen2.5-Coder-32B (98.84%) but moderate success elsewhere. Preprocessor carriers consistently outperform identifier substitution, suggesting over-reliance on non-executable boundary tokens. GPT-5-mini exhibits uniformly low ASR (10–28%), indicating broader distributional invariance than other models. Dead-branch carriers are effective across multiple models, indicating limited dead-code resolution in current detectors.
Model-specific patterns. Most models show high vulnerability to preprocessor carriers; Llama3.1-8B and Qwen2.5-Coder-32B are near-ceiling (85–99%). CodeAstra presents a distinct profile: identifier substitution transfers poorly (4.72%) while comment and preprocessor carriers remain effective (42–79%); on-target optimization reverses this (§5.4). Identifier substitution ASRs are computed over the full $\mathrm{TP}_{\mathrm{clean}}$ set, with compilation failures (9.1%; Table 4) counted as unsuccessful attacks (conservative estimate). StarCoder2-15B exhibits a single-carrier dependency: macro alone accounts for nearly all transfer evasion (86.83%), confirming that one vulnerable surface suffices to undermine overall robustness.
@@
5.3 Defense Evaluation
We evaluate three defenses of increasing strength, spanning inference-time preprocessing, inference-time ensembling, and training-time intervention.
@@ 5.3 Defense Evaluation
Randomized smoothing. Adapting SmoothLLM Robey et al. (2024) to vulnerability detection, we generate K=5K{=}5 variants per input via random identifier renaming (up to 2 renames) and take majority vote. We evaluate this on the identifier substitution carrier under on-target GCG. ASR reductions are modest across all three evaluated models: Llama3.1-8B ($43.49\%\to 34.57\%$, 8.9-8.9 pp), CodeAstra ($85.41\%\to 82.71\%$, 2.7-2.7 pp), StarCoder2-15B ($27\%\to 14.63\%, 1.6-1.6$ pp). This confirms that GCG-optimized adversarial identifiers are robust to small random perturbations of the input space.
Adversarial fine-tuning. To evaluate a training-time defense, we LoRA fine-tune Llama3.1-8B on carrier-diverse adversarial examples spanning all five carriers and re-optimize new universal strings (with a held-out seed) on the hardened model. Comment and preprocessor carriers drop sharply (comment: $42.01\%\to 0.16\%$; ifdef: $99.03\%\to 0.00\%$; macro: $65.46\%\to 16.91\%$), but identifier substitution ASR increases ($43.49\%\to 89.56\%$) and dead-branch remains high ($86.66\%\to 68.20\%$). Overall CR improves from 0.13% to 3.88%, while clean TPR drops from 31.04% to 24.72% (6.3-6.3 pp). Adversarial training thus closes specific carrier surfaces while potentially exposing others, and introduces a measurable robustness-accuracy trade-off.
Cross-defense summary. No single defense is sufficient. Sanitization eliminates three carrier surfaces but leaves two intact and distorts the decision boundary. Smoothing yields marginal gains against optimized perturbations. Adversarial training provides the largest CR improvement but transfers unevenly across carriers. Effective mitigation likely requires combining complementary defenses across inference and training time; per-model breakdowns are in Appendix D.
@@ 5.4 Transfer Attacks Succeed on Black-Box APIs; On-Target Reveals Complex Interactions
Table 2 presents transfer results where universal strings optimized on Qwen2.5-Coder-14B are applied unchanged to all targets. This models a low-barrier adversary: optimization is performed once, and resulting strings are reused across vulnerable functions and deployed detectors. Across black-box API and cost-limited gray-box targets, transfer achieves 44.04%–99.38% union $\mathrm{ASR}_{\mathrm{cond}}$, establishing that API opacity alone does not prevent evasion under syntax-preserving perturbations. However, transferability is not uniform: GPT-5-mini exhibits lower success (44.04%) than other models (87–99%). To test whether this reflects genuine robustness or insufficient surrogate capacity, we applied identifier substitution strings optimized on-target on Qwen2.5-Coder-32B (a stronger surrogate) to GPT-5-mini; ASR increases only marginally (10.62% $\to$ 13.4%), suggesting that the lower transferability is not an artifact of weak surrogate transfer.
@@
On-target optimization reveals divergent vulnerability profiles. While transfer attacks establish a practical lower bound on exploitability, they may underestimate the true vulnerability of a model. To quantify the gap between low-cost transfer and worst-case security risk, we perform on-target universal GCG forLlama3.1-8B, CodeAstra, and StarCoder2-15B (white-box models); full per-carrier breakdown in Table 7, Appendix C.2). StarCoder2-15B shows consistent gains under on-target optimization across most carriers (e.g., comment: 4.00% $\to$ 69.59%), reducing CR from 12.28% to 9.61%.
@@
For Llama3.1-8B, on-target optimization improves identifier substitution ($34.66\%\to 43.49\%$) and dead-branch ($59.79\%\to 86.66\%$) but degrades comment ($64.50\%\to 42.01\%$) and macro ($96.33\%\to 65.46\%$). Universal strings optimized on a surrogate can outperform target-specific optimization for certain carriers, likely because cross-model training discovers more transferable boundary-token sensitivities. CodeAstra exhibits a clear asymmetry. Transfer achieves 4.72% identifier success but 42.59% for ifdef. On-target optimization inverts this: identifier reaches 85.41% while ifdef collapses (16.62%).
@@
We hypothesize two mechanisms. For identifier substitution, CodeAstra’s loss landscape provides smooth gradients: GCG efficiently discovers variable-name perturbations that flip predictions. For #ifdef, on-target GCG likely encounters optimization difficulty, converging to poor local minima that fail to exploit the vulnerabilities that surrogate-optimized strings discovered through multi-model training. This suggests gradient-based optimization success depends not only on model vulnerabilities but also on the optimizability of the attack surface. A related observation on Qwen2.5-Coder-14B (identifier substitution converges smoothly while comment insertion oscillates; Appendix C.3) provides indirect support; direct loss-landscape analysis on CodeAstra #ifdef is left to future work.
@@
Qwen2.5-Coder-32B: Surrogate capacity and on-target upper bound. For Qwen2.5-Coder-32B (cost-limited gray-box), we ablate identifier substitution across surrogate capacities and on-target optimization (Table 8, Appendix C.2). Transferability is non-zero even for small surrogates (0.5B/1.5B:  6-7%), but increases with capacity (14B: 26.94%), indicating a representation-capacity threshold. On-target universal GCG achieves 92.60%, providing a clear upper bound when gradients are available. This gap (26.94% $\rightarrow$ 92.60%) quantifies the security improvement from restricting adversarial access: surrogate transfer yields a conservative lower bound under practical compute constraints, while on-target optimization demonstrates worst-case exploitability.
@@
5.5 Non-GCG Baselines